Paul has a son named David. The sum of their ages is 84. After three years, Paul will be twice as old as David. What are their ages now?
Answer by jorel1380(3719)

Let David be d. Then Paul is 84-d. So:
2(d+3)=84-d+3
2d+6=87-d
3d=81
d=27
David is 27; Paul is 57. ☺☺☺☺